# MedAide: Towards an Omni Medical Aide via Specialized LLM-based Multi-Agent Collaboration

### v2

@@ Abstract
Large Language Model (LLM)-driven interactive systems currently show potential promise in healthcare domains. Despite their remarkable capabilities, LLMs typically lack personalized recommendations and diagnosis analysis in sophisticated medical applications, causing hallucinations and performance bottlenecks. To address these challenges, this paper proposes MedAide, an LLM-based omni medical multi-agent collaboration framework for specialized healthcare services. Specifically, MedAide first performs query rewriting through retrieval-augmented generation to accomplish accurate medical intent understanding. Immediately, we devise a contextual encoder to obtain intent prototype embeddings, which are used to recognize fine-grained intents by similarity matching. According to the intent relevance, the activated agents collaborate effectively to provide integrated decision analysis. Extensive experiments are conducted on four medical benchmarks with composite intents. Experimental results from automated metrics and expert doctor evaluations show that MedAide outperforms current LLMs and improves their medical proficiency and strategic reasoning.
MedAide: Towards an Omni Medical Aide via Specialized
LLM-based Multi-Agent Collaboration
@@ 1 Introduction
Developing the goal-oriented conversation system Lee et al. (2024); Cheng et al. (2023) has received increasing attention in recent years. Advanced Large Language Models (LLMs) OpenAI (2022); Achiam et al. (2023); OpenAI (2024) have shown exceptional generalization capabilities in general scenarios for multiple purposes within human-machine interactions, such as negotiation He et al. (2018) and persuasion Wang et al. (2019). In this context, LLM-centered interactive medical assistants Bao et al. (2023); Yang et al. (2024a, 2023b); Shi (2023); Chen et al. (2023b) have become research hotspots that promise to improve diagnosis efficiency and promote service automation. Previous attempts infuse LLMs with healthcare-specific knowledge through relevant corpus construction (e.g., knowledge databases Li et al. (2023b)) and multi-stage training procedures (e.g., supervised fine-tuning Xiong et al. (2023)). Although these strategies stimulate the models’ comprehension of medically relevant intents, there are still bottlenecks when faced with real-world applications that require sophisticated reasoning.
@@
Given the mimicking of learned feedback in human behaviors Du et al. (2023); Park et al. (2023), automated medical agent construction promises to enhance LLMs’ instruction following and logical profiling capabilities Lee et al. (2024). The collaboration among multiple agents to handle different patient inquiries and symptomatic caseloads facilitates accurate dialog goal fulfillment while accounting for individual differences Fan et al. (2024); Li et al. (2024). Despite significant advancements, current efforts focus primarily on medical education training Wei et al. (2024) or selective question-answering Tang et al. (2024), lacking understanding of complicated intents behind user queries and giving hierarchical therapeutic recommendations.
To address these issues, this paper proposes an LLM-based omni medical agent framework for real-world medical applications, called MedAide. The core philosophy of MedAide lies in first complementing and decomposing multi-dimensional medical intents through query rewriting to enhance the model’s dissection of composite medical queries. Then, we introduce a contextual encoder to learn intent prototype embeddings from multifaceted instructions, which is utilized to specify fine-grained target intents via the similarity matching lemma. In this case, different paramedical agents are correspondingly activated to provide personalized determinations with specialized medical expertise. Ultimately, we formulate a decision analysis module with Chain-of-Thought (CoT) properties to summarize the responses of activated agents in a faithful manner to make integrated decisions. The main contributions are summarized as follows:
  • To the best of our knowledge, we are the first to propose the omni multi-agent collaboration framework for real-world scenarios with composite healthcare intents, which shows potential for advancing interactive systems for personalized healthcare.
  • Our MedAide effectively improves the strategic reasoning of LLMs under complicated dialog goals through feedback and collaboration of specialized paramedical agents.
  • Extensive experiments on seven medical benchmarks spanning 17 types of rich intents prove the effectiveness of MedAide. As a plug-and-play framework, MedAide can be readily combined with current LLMs and provides competitive improvements.
@@ 2.1 LLMs in Healthcare Domains
The Large Language Models (LLMs) exemplified by ChatGPT (OpenAI, 2022) and GPT-4 Achiam et al. (2023) demonstrate excellent performance in multidisciplinary applications. Although current LLMs Bai (2023); Yang et al. (2023a); AI (2024) with certain medical knowledge benefit from large-scale corpus support, they lack specialized medical proficiency and have significant performance bottlenecks in domain-specific scenarios. Recently, several attempts Chen et al. (2023b, a); Xu (2023); Bao et al. (2023); Yang et al. (2024a) have begun to build medically customized LLM assistants to fulfill the diagnostic and consultative demands. For instance, HuatuoGPT series Zhang et al. (2023); Chen et al. (2023a) have shown promising results in bridging generic-medical knowledge gaps by absorbing real doctor-patient conversations. ZhongJing series Yang et al. (2023b); Shi (2023) improve Chinese medical capabilities by introducing expert feedback and multi-round medical instructions. In addition, PediatricsGPT Yang et al. (2024a) proposes a systematic training framework to construct interactive healthcare systems for pediatric specialists and medical generalists. Unlike previous studies, our framework aims to more fully recognize medical intents and refine the models’ reasoning abilities in complex scenarios through the LLM-based multi-agent collaboration.
@@ 2.2 LLM-based Multi-agent Collaboration
With the focus of researchers on sophisticated goal-oriented dialog generation Wu et al. (2023); Fan et al. (2024); Jiang et al. (2024), the inherent dilemmas of hallucinatory responses and weak comprehension Chen et al. (2024); Yang et al. (2024b) in LLMs have been gradually exposed. In this context, LLM-based automated agents are proposed to provide effective perception and decision-making skilfulness by incorporating external tools and databases Cai et al. (2023); Li et al. (2023a). By mimicking human behavioral logic, multiple agents perform feedback and collaboration to enhance diverse intent understanding tasks, including educational training Lee et al. (2024); Wei et al. (2024), emotional comfort Cheng et al. (2023), and workflow integration Hong et al. (2023). For example, MEDCO Wei et al. (2024) enables LLMs to simulate patients and doctors to enhance the practice performance of virtual students in interactive environments. MedAgents Tang et al. (2024) improves the performance of medical assistants in zero-shot settings through the role-playing strategy. In comparison, the proposed MedAide focuses more on mining profound healthcare intents and moving towards robust healthcare practices.

### v3

# MedAide: Information Fusion and Anatomy of Medical Intents via LLM-based Agent Collaboration

@@ Abstract
In healthcare intelligence, the ability to fuse heterogeneous, multi-intent information from diverse clinical sources is fundamental to building reliable decision-making systems. Large Language Model (LLM)-driven information interaction systems currently showing potential promise in the healthcare domain. Nevertheless, they often suffer from information redundancy and coupling when dealing with complex medical intents, leading to severe hallucinations and performance bottlenecks. To this end, we propose MedAide, an LLM-based medical multi-agent collaboration framework designed to enable intent-aware information fusion and coordinated reasoning across specialized healthcare domains. Specifically, we introduce a regularization-guided module that combines syntactic constraints with retrieval-augmented generation to decompose complex queries into structured representations, facilitating fine-grained clinical information fusion and intent resolution. Additionally, a dynamic intent prototype matching module is proposed to utilize dynamic prototype representation with a semantic similarity matching mechanism to achieve adaptive recognition and updating of the agent’s intent in multi-round healthcare dialogues. Ultimately, we design a rotation agent collaboration mechanism that introduces dynamic role rotation and decision-level information fusion across specialized medical agents. Extensive experiments are conducted on four medical benchmarks with composite intents. Experimental results from automated metrics and expert doctor evaluations show that MedAide outperforms current LLMs and improves their medical proficiency and strategic reasoning.
[…]
@@ 1 Introduction
The emergence of information fusion technologies has revolutionized healthcare systems by enabling the integration of multi-source medical data for enhanced decision-making capabilities [33]. Contemporary multi-source information fusion approaches have proven crucial for processing diverse healthcare data modalities, enabling comprehensive patient assessment and personalized treatment strategies [23]. In the healthcare domain, the integration of artificial intelligence with information fusion techniques has demonstrated significant potential for smart healthcare applications [3]. The advancement of multimodal medical signal fusion has played a crucial role in smart healthcare systems, enabling real-time analysis and providing medical professionals with enhanced insights for disease detection and prediction accuracy [27]. Furthermore, multi-modal lifelog data fusion approaches have shown promising results in improving human activity recognition and health monitoring applications [15]. Advanced deep learning technologies [29, 30, 35, 20, 40, 37, 26, 25] have shown exceptional generalization capabilities in general scenarios for multiple purposes within human-machine interactions. Developing goal-oriented conversation systems [8] has received increasing attention in recent years, particularly in medical contexts where information fusion enables sophisticated reasoning across heterogeneous data sources. In this context, LLM-centered interactive medical assistants [40, 41, 7] have become research hotspots that promise to improve diagnosis efficiency and promote service automation. The advancement of multimodal medical signal fusion has played a crucial role in smart healthcare systems, enabling real-time analysis and providing medical professionals with enhanced insights for disease detection and prediction accuracy [27]. Previous attempts have infused LLMs with healthcare-specific knowledge through relevant corpus construction and multi-stage training procedures . However, traditional approaches often overlook the potential of multi-source information fusion for achieving a comprehensive understanding of complex medical scenarios [15]. Although these strategies enhance the models’ comprehension of medically relevant intents, bottlenecks remain when faced with real-world applications that require sophisticated reasoning and accurate feedback, particularly in scenarios that demand the integration of diverse information sources for optimal clinical decision-making.
@@
Given the mimicking of learned feedback in human behaviors, automated medical agent construction promises to enhance LLMs’ instruction following and logical profiling capabilities. The collaboration among multiple agents to handle different patient inquiries and symptomatic caseloads facilitates accurate dialog goal fulfillment while accounting for individual differences [12]. Despite significant advancements, current efforts focus primarily on medical education training [36] or selective question-answering [34], lacking a comprehensive understanding when dealing with the sophisticated intents behind user queries in real-world diagnosis and treatment scenarios. In addition, the agents in most collaborative frameworks are usually familiar with only limited intents, making it difficult to provide systematic recommendations.
Figure 1: Comparison between our framework and previous methods. Among these works, MedAide is the only one with intent awareness and comprehensive analysis capabilities.
To address these issues, we propose MedAide, an LLM-based medical multi-agent collaboration framework designed to enable intent-aware information fusion and coordinated reasoning across specialized healthcare domains. There are three core contributions in MedAide based on the tailored components. Specifically, we propose a regularization-guided information extraction module, which fuses syntactic constraints with retrieval-augmented generation to transform compositional medical queries into structured semantic representations. This structured decomposition enhances the granularity and reliability of clinical knowledge extraction, facilitating downstream intent interpretation. In addition, a dynamic intent prototype matching mechanism is introduced to adaptively recognize and update agent intent through semantic similarity-based alignment during multi-turn healthcare dialogues. Ultimately, we design a rotation-based agent collaboration mechanism that enables dynamic role-switching and decision-level information fusion across agents, further ensuring coherent and globally optimized decision-making. Experimental results on four medical benchmarks with composite intents, evaluated by both automated metrics and expert physicians, demonstrate that MedAide significantly improves medical reasoning accuracy, intent alignment, and agent coordination over strong LLM baselines.
Figure 2: Illustration of the MedAide framework. MedAide consists of three major phases: (i) Regularization-guided Information Extraction (RIE) employs syntactic parsing and retrieval-augmented generation to reformulate complex medical queries and extract clinical elements. (ii) Dynamic Intent Prototype Matching (IPM) utilizes a BioBERT-based encoder to generate intent embeddings for precise classification across 17 medical intent categories. (iii) Rotation Agent Collaboration (RAC) implements a polling-based mechanism where specialized agents (pre-diagnosis, diagnosis, medicament, post-diagnosis) dynamically rotate as main contacts, coordinating multi-source information fusion through cross-agent knowledge integration.
@@ 2 Related Work
2.1 LLM-enhanced Information Fusion
The emergence of Large Language Models (LLMs) has introduced revolutionary paradigms for information fusion, transforming the integration and processing of heterogeneous data sources. Recent advancements demonstrate that LLMs serve as sophisticated fusion engines capable of harmonizing diverse information types and enabling semantic understanding across different data domains [28]. Contemporary research has demonstrated several key approaches to LLM-based information fusion, including context-based fine-tuning frameworks that leverage information fusion principles to enhance model adaptability while preserving pre-trained knowledge [28]. Novel techniques such as knowledge distillation and attention-based fusion mechanisms enable seamless integration of heterogeneous information sources into LLMs, representing significant advancement in intelligent data processing capabilities [43]. Various application domains have particularly benefited from LLM-enhanced information fusion approaches, where LLMs demonstrate exceptional capabilities in understanding, reasoning, and generation, introducing transformative paradigms for integrating artificial intelligence into complex decision-making systems [39]. These developments encompass fusion strategies operating at various architectural levels, from feature-level integration to decision-level fusion, with emerging hybrid approaches that leverage the representational power of LLMs to bridge semantic gaps effectively.
@@ 2.2 LLMs in Healthcare Domains
Large Language Models (LLMs), exemplified by ChatGPT [29], demonstrate excellent performance in multidisciplinary applications. Although current LLMs [1, 2] with certain medical knowledge benefit from large-scale corpus support, they lack specialized medical proficiency and have significant performance bottlenecks in domain-specific scenarios. Recently, several attempts [7, 6, 40] have begun to build medically customized LLM assistants to fulfill the diagnostic and consultative demands. For instance, HuatuoGPT series [6] have shown promising results in bridging generic-medical knowledge gaps by absorbing real doctor-patient conversations. ZhongJing [41] improves Chinese medical capabilities by introducing expert feedback and multi-round medical instructions. In addition, PediatricsGPT [40] proposes a systematic training framework to construct interactive healthcare systems for pediatric specialists and medical generalists. Unlike previous studies, our framework aims to more fully recognize medical intents and refine the models’ reasoning abilities in complex scenarios through the LLM-based multi-agent collaboration.
@@ 2.3 LLM-based Multi-agent Collaboration
With the focus of researchers on sophisticated goal-oriented dialog generation [12], the inherent dilemmas of hallucinatory responses and weak comprehension in LLMs have been gradually exposed. In this context, LLM-based automated agents are proposed to provide effective perception and decision-making skilfulness by incorporating external tools and databases [22]. By mimicking human behavioral logic, multiple agents perform feedback and collaboration to enhance diverse intent understanding tasks, including educational training [36] and emotional comfort [8]. For example, MEDCO [36] enables LLMs to simulate interactions between patients and doctors, enhancing the practice performance of virtual students in interactive environments. MedAgents [34] improves the performance of medical assistants in zero-shot settings through the role-playing strategy. MDAgents [18] flexibly select the corresponding processing method according to the complexity of the inquiry. A detailed comparison between our framework and previous methods is provided in Figure  1. In comparison, the proposed MedAide focuses more on mining profound healthcare intents and moving towards robust healthcare practices.